a) (2 pts) Suppose that setting up a call requires reserving 3 link segments. Each segment is available with probability 0.3. What is the probability that a call request can be completed?

b) (2pts) Explain where the following fits in the OSI reference model: In WWII, a teleprinter transmits a message over radio (telex-on-radio, TOR).
Answer #1

`Hey,

Note: Brother if you have any queries related the answer please do comment. I would be very happy to resolve all your queries.

a)

For a call to be completed, every segment in the circuit must be free. Thus, assuming that each switch and link has a probability p of being available, and assuming that this probability is independent of all others,

P[every circuit free]= P[1st free]⋅P[2nd free]⋅...⋅ P[Nth free]= p^n

So, given p=0.3 and n=3

SO,

P=(0.3)^3=0.027

Note: Brother According to Chegg's policy we are only allowed to answer first part if there are many. So, I request you to post other part as separate posts

Kindly revert for any queries

Thanks.